# MATHGLANCE: Multimodal Large Language Models Do Not Know Where to Look in Mathematical Diagrams

### v1

@@ 2 MATHGLANCE
Table 1: Key statistics of MATHGLANCE are summarized in Tab. 3(a), and the subject-task distribution is illustrated in Fig. 3(b).
@@
2.1 Overview
Existing mathematical visual reasoning benchmarks often conflate perception with higher-level reasoning tasks, such as numeric calculations and proof generation. We introduce MATHGLANCE, a novel benchmark designed to evaluate MLLMs’ mathematical perception-demanding abilities through both quantitative and qualitative analysis across coarse-to-fine granularity levels.
Data Composition. Detailed statistics for data composition are presented in Tab. 3(b). To comprehensively assess the perceptual abilities of MLLMs in mathematical contexts, our evaluation images cover plane geometry (66%), solid geometry (20%), and graphs (14%), including lines, bars, and pie charts. To facilitate MLLM evaluation, we formulate all tasks—except for bounding box grounding (free-form)—as multiple-choice or true/false question-answering problems. In total, we contribute 1,609 questions and 1,198 unique images, ensuring an even distribution across the different perception tasks.
Categorization. The benchmark encompasses shape classification, object counting, relationship identification, and object grounding. Fig. 3 shows example illustrations of plane geometry (see §B for additional examples). Key features are as follows:
  • Shape classification is a classic vision task where the model identifies an object’s class based on its attributes, i.e., vertices, material, color, and size. Our dataset includes a diverse set of geometric categories, comprising 16 basic shapes for plane geometry, 3 CLEVR-defined [24] objects for solid geometry, and 5 graphical elements as defined in FigureQA [25].
These four tasks are designed to isolate and assess fundamental perception capabilities, distinguishing low-level visual recognition from high-level reasoning.
Figure 4: The synthetic construction process for plane geometry. We synthesize geometric figures by randomly sampling elements from the geometric shape pool and relationship pool, ensuring consistency through a verifier that enforces logical constraints based on manually designed rules, fundamental mathematical principles, and prerequisite points. All visual elements are structured and saved in JSON format. Images are rendered using the Matplotlib package, and corresponding Q&A pairs are generated using a template-based pipeline.
2.2 Dataset Construction
2.2.1 Synthetic Data Engine for Plane Geometry
We design a data generation engine to synthesize real-world geometric images along with structured JSON annotations. This pipeline enables controlled shape generation, relationship modeling, and visual attribute assignment, ensuring diverse and well-balanced datasets across perception tasks.
Structured Annotations for Geometric Primitives. Fig. 4 describes the entire generation process. Inspired by AlphaGeometry [43], we use geometric clauses as fundamental units to construct complex plan geometric figures. A geometric clause is a formalized description of basic geometric objects and mathematical relationships, along with their properties or attibutes, a.k.a. prerequisite points. We first construct two geometry substrate pools, one containing 16 different geometric shapes (i.e., scalene triangle, isosceles triangle, square, rectangle, parallelogram, isosceles trapezoid, right trapezoid, pentagon, circle $\dots$ ellipse and segment), and the other defining 10 mathematical relationships (i.e., on, intersection, parallel, perpendicular, tangent, incircle, $\dots$ reflection and angle bisector). We then randomly sample one or more substrates from these pools and pass them through a verifier, which ensures that logically paired shapes and relationships are preserved in the construction of valid geometric images. The verifier makes decisions based on either manually designed rules, fundamental mathematical knowledge, or prerequisite points. For example, parallel lines cannot intersect, making an intersection clause invalid. Similarly, angle trisection requires a predefined angle, which cannot be constructed without enough prerequisite points for angle consturction. In general, the chosen relationships are enforced by introducing additional shapes into the figure, ensuring that the specified relationships are accurately maintained and geometrically consistent throughout the construction process. We finally save the outputs as structured JSON annotations for image rendering using the Matplotlib package [22] and generating question-answer pairs via template-based pipelines.
Image Generation. The JSON annotations define foreground and background styles, i.e., colors sampled from a monochromatic palette, line width, and font size, as well as object shape information (class names), attributes (vertices labeled with random letters), bounding box locations, and mathematical relationships. Notably, spatial relationships are generated based on the bounding box locations of two objects (e.g., top-left, top-right, bottom-left, bottom-right). Similar to the image rendering process in  [43], we translate geometric clauses into visual representations using Python code. This process first determines the coordinates of each point defined in the clause for basic shape visualization, and then new primitives are generated based on the specified relationships. Additionally, we apply image augmentations to increase task difficulty, e.g., adding Gaussian noise, drawing irregular scribbles, and incorporating wedge-shaped symbols for congruent angles or auxiliary lines (see Figs. 9 and 11 in the Appendix).
Question-Answer Generation. Based on the structured annotations, we employ a template-based pipeline to generate multiple-choice and true-false questions across four perception tasks. For example, in constructing Q&A pairs for shape classification, we first randomly select an object and its associated attributes from the given figure. We then formulate a question by filling placeholders in our carefully designed templates. For true-false questions, the ground truth answer is directly derived from the JSON annotations. For multiple-choice questions, we generate plausible distractors to challenge the model’s geometric perception, combining them with the correct answer into a set of four choices. Distractors are selected from geometric candidate pools, ranging from visually similar (e.g., equilateral vs. isosceles triangles) to visually dissimilar pairs (e.g., equilateral triangle vs. circle). Further details on the Q&A generation templates are provided in Tab. 7 of the Appendix.
2.2.2 Reformat Dataset for Solid Geometry & Graphs
We leverage the public datasets CLEVR [24] and FigureQA [25] to construct benchmarks for solid geometry and graphs, respectively. Since their original annotation formats are incompatible with our template-based Q&A generation pipeline, we reformat them into structured JSON, consistent with the format used in our synthetic data process. Once reformatted, we apply the same Q&A generation strategy as used for plane geometry. See §D for details.
Table 2: Performance comparison of different MLLMs on MATHGLANCE across plane geometry, solid geometry, and graphs. cls, cnt, grd, and rlat represent different question categories: shape classification, object counting, object grounding, and relationship identification, respectively. all indicates the overall accuracy, calculated as the ratio of correctly answered questions to the total number of questions in the benchmark, while Avg. denotes the average all score across all subjects.
Model Size Avg. Plane Geometry Soild Geometry Graphs
all cls cnt grd rlat all cls cnt grd rlat all cls cnt grd rlat
Open-Source Generic MLLMs
Closed-Source Generic MLLMs
Open-Source Mathematical MLLMs
MultiMath [38] 7B 42.1 31.2 44.0 30.4 1.07 53.0 46.7 81.4 53.6 4.7 33.8 48.6 79.7 57.4 3.2 33.8
SVE-Math-DeepSeek [55] 7B 46.6 35.4 52.4 36.0 3.56 51.0 49.4 77.9 62.9 1.5 41.3 55.1 81.2 75.9 1.6 69.7
SVE-Math-DeepSeek+ (ours) 7B 68.4 84.6 75.8 88.4 82.9 97.5 54.1 85.3 65.8 20.3 45.0 60.7 85.1 78.4 1.6 75.7
2.3 Perception-oriented Training Dataset (GeoPeP)
Evaluation results on MATHGLANCE reveal that both open-source and closed-source MLLMs struggle to identify relevant visual regions in symbolic and abstract mathematical diagrams, despite their strong performance in theoretical reasoning and numerical computation. This contrasts with human cognitive abilities, where low-level perceptual tasks are typically solved rapidly compared to high-level reasoning tasks. To overcome such perceptual limitations, we construct a perception-oriented training dataset, GeoPeP, which enhances geometric understanding and visual grounding, ultimately benefiting complex reasoning. GeoPeP, comprising 200K high-quality image-text pairs in both diagram-caption and instruction-conversation formats, provides detailed geometric primitive information, including shapes, vertices, relationships, and fine-grained bounding box coordinates.
[…]
Overall, GeoPeP contributes to MLLM training by: (1) providing clear object attributes and their relationships, akin to graph nodes and edges, in training QA pairs; (2) offering fine-grained bounding box coordinates of elements, enabling models to systematically learn spatial awareness; and (3) integrating with reasoning-based CoT mathematical visual datasets during self-fine-tuning stage, allowing models to both perceive and reason accurately.
3 Experiments
In this section, we first describe the experimental setup (§3.1). We then present a comprehensive evaluation of 15 recent MLLMs (§3.2), demonstrating that while humans achieve high accuracy, MATHGLANCE remains challenging for existing models. Finally, we conduct comprehensive ablations and analyses, including detailed comparisons with math-specific MLLMs on both reasoning and perception benchmarks; investigation of key factors affecting MLLMs’ perception during inference; and evaluation of how GeoPeP improves both perception and reasoning (§3.3).
3.1 Experimental Setup
Generic and Mathematical MLLMs. We evaluate 15 MLLMs on MATHGLANCE across plane geometry, solid geometry, and graphs, including closed-source generic models such as GPT-4o and GPT-o1, as well as open-source models like LLaVA-v1.5 [32], mPLUG-Owl3 [48], InternLM-XComposer2 [16], Qwen2VL [46], Qwen2.5VL [6], DeepSeek-VL2 [47], InternVL2 [11], and InternVL2.5 [10]. Additionally, we assess math-specific MLLMs, including SVE-Math-DeepSeek [55], Math-LLaVA [41], G-LLaVA [18], and MultiMath [38]. This comprehensive evaluation provides insights into the diagram perception capabilities of state-of-the-art multimodal models.
Implementation Details. For open- and closed-source generic MLLMs, we follow the official inference settings, including temperature, number of beams, and maximum token length. For open-source mathematical LLMs, we adopt the standard configurations from SVE-Math-DeepSeek, setting the temperature to 0, the number of beams to 1, and the maximum token length to 1024. Notably, MultiMath [38] uses a temperature of 0.2; therefore, for a fair comparison, we rerun the inference with the temperature set to 0. We retain the original 512$\times$512 image resolution throughout all experiments. Choices are extracted using predefined rules tailored to each MLLM’s output format.
3.2 Main Results
The performance comparison of current MLLMs on plane geometry, solid geometry [24], and graphs [25] is summarized in Table 2. To assess models’ geometric diagram perception abilities, we evaluate their performance across shape classification, object counting, relationship identification, and object grounding. Key findings reveal that both generic and mathematical MLLMs face significant challenges in geometric perception tasks.
Generic MLLMs. General-purpose models trained on diverse datasets, including tables, charts, and documents [25, 9, 51], as well as visual grounding datasets [49, 40], still perform poorly on mathematical diagram perception. In particular, their performance on plane geometry remains low, with most models scoring below 45% on average. For solid geometry and graphs, general-purpose models significantly outperform mathematical MLLMs, due to their exposure to large-scale FigureQA [25], CLEVR [24], and various chart understanding datasets [51]. However, they still fail in fine-grained box-level tasks, with most models achieving 0 accuracy, and lag significantly behind human-level perception. These evaluation results highlight the limitations of generic MLLMs in mathematical diagram perception. Despite being trained on large-scale and diverse natural image datasets (e.g., over 2T tokens for Qwen2.5-VL and DeepSeek-VL2), these models struggle with the abstract nature of mathematical diagrams, which are defined by precise geometric structures and symbolic notations. Unlike natural images—where models can exploit semantic priors (e.g., people typically appear on the ground, birds in the sky)—mathematical diagrams lack such contextual cues. Their symbolic and structured nature requires genuine visual understanding rather than superficial pattern recognition. Without this, models fail to identify where to look, leading to poor perception performance. Notably, scaling up model size is neither an optimal nor an effective solution, as it provides only marginal gains in perception compared to reasoning benchmarks. For instance, increasing Qwen2VL from 7B to 72B improves top-1 accuracy by 22.3% on MathVista but only 8.3% on MATHGLANCE.
Mathematical MLLMs. Models designed for mathematical reasoning, such as MultiMath, Math-LLaVA, and SVE-Math-DeepSeek, exhibit strong reasoning performance on benchmarks like MathVerse [53], MathVista [34], and GeoQA [18]. They struggle with fundamental geometric perception, particularly fine-grained object localization, achieving near-zero accuracy at the default IoU threshold of 0.65. Since SVE-Math-DeepSeek incorporates a geometric visual encoder trained with detection and boundary segmentation losses, it outperforms other mathematical MLLMs on our benchmark across both cross-level and fine-level perception tasks. Furthermore, as it is not exposed to solid geometry diagrams or graphical images during MLLM training stages, it provides a more controlled experimental setting for analyzing perception capabilities compared to general-purpose models. Therefore, we select SVE-Math-DeepSeek as our base model to explore enhancements in both perception and reasoning. The comparison between SVE-Math-DeepSeek and other mathematical MLLMs is shown in Tab. 3. SVE-Math-DeepSeek trained on GeoPeP significantly improves solid geometry perception. Even without direct training on solid geometry and graphs, our model outperforms others due to its enhanced ability to discriminate relationships and understand spatial configurations. Low-level diagram structure understanding is essential for high-level reasoning skills, contributing to $\sim+4\%$ across three mathematical reasoning benchmarks. We leave the exploration of structure-aware training samples for solid geometry and graphs to future work.
Table 3: Performance comparison of mathematical MLLMs on both math perception and reasoning benchmarks.
Model MATHGLANCE MathVerse MathVista GeoQA
Plane Soild Graphs
Math-LLaVA [41] 27.9 44.8 47.3 20.1 46.6 60.7
G-LLaVA [18] 25.6 31.3 33.9 17.8 25.6 64.2
MultiMath [38] 31.2 45.7 48.6 25.9 49.3 74.1
SVE-Math-DeepSeek [55] 35.4 49.4 55.1 24.3 48.7 72.8
SVE-Math-DeepSeek+ 84.6 54.1 60.7 28.1 51.3 76.2
Figure 5: We evaluate five key factors affecting the perception ability of MLLMs: the number of obejcts (# obejct), visual quality, visual and textual distractors, and Chain-of-Thought (CoT) reasoning strategies (see Fig. 6 for details).
Figure 6: Analyzing perceptual errors: CoT vs. direct reasoning.
Figure 7: Response comparisons between SVE-Math-DeepSeek $\!\!{}^{+}$, SVE-Math-DeepSeek, and InterVL2.5.
3.3 Ablation Study
3.3.1 Key Factors Influencing Diagram Perception
We examine five key factors: the number of objects, object quality, visual distractors, text distractors, and Chain-of-Thought (CoT) responses. While those factors are by no means exhaustive, they aim to illuminate some of the fundamental perceptual limitations of current MLLMs, providing insights for practical applications and future improvements. The evaluation results on plane geometry from MATHGLANCE across nine MLLMs are presented in Fig.5, and response error patterns are shown in Fig.6.
The Number of Objects. The complexity of visual perception tasks increases with the number of objects in an image. We observe that as the number of geometric shapes grows, models struggle with accurate object localization, shape classification, and counting. Fig. 1 presents the overall results across all perception tasks for eight MLLMs and our method. When scaling the number of objects from 1 to $N$ ($N<5$), the Qwen-series model’s performance drops by approximately 12%, the DeepSeek-series by 18%, and GPT-4o by 15%. Our method and its baseline show declines of 6% and 13%.
MATHGLANCE Noise $\sim\mathcal{N}({0},{0.3})$ Distractors
w/o w/Δ↑↓ w/o w/Δ↑↓
Qwen2VL-7B 29.8 33.9↑4.1 29.8 26.2↓3.6
Qwen2.5VL-7B 33.9 32.1↓1.8 33.9 31.0↓2.9
DeepSeek-VL2-Tiny 27.9 23.8↓4.1 27.9 22.6↓5.3
DeepSeek-VL2-Small 27.9 37.5↑9.6 27.9 25.9↓2.0
LLaVa-1.5-13B 26.8 31.6↑4.8 26.8 29.1↑2.3
InternVL2.5-8B 25.6 20.8↓4.8 25.6 19.6↓6.0
GPT-4o 32.7 30.9↓1.8 32.7 28.6↓4.1
SVE-Math-DeepSeek-7B 32.8 28.0↓4.8 32.8 29.2↓3.6
SVE-Math-DeepSeek +-7B 80.7 78.3↓2.4 80.7 80.3↓0.4
Visual Distractors. We draw visual distractors—irregular scribbles, wedge-shaped angle markers, and auxiliary lines—to evaluate MLLMs’ ability to focus on relevant geometric elements (see Fig.9). Despite prompting models to ignore these distractors, most are still negatively affected. Close-up results are shown in Tab. 4. All evaluated MLLMs show a performance drop of 2–10% under visual distractors, except LLaVA-1.5-13B, which improves by 2.3%. Our model remains robust, showing minimal sensitivity to distractors and better visual focus on geometric primitives.
Object Quality. For visual fidelity analysis, we apply Gaussian noise to degrade object quality. As shown in Tab. 4, with a default variance of 0.3, most models show performance degradation, while Qwen2VL, DeepSeek-VL2-Small, and LLaVA-1.5-13B unexpectedly improve. This may stem from their exposure to visually degraded data during training, making noisy images more aligned with their learned distribution—especially in the absence of clean geometric diagrams in the training corpus. To enable a more comprehensive evaluation and mitigate data bias, we introduce large Gaussian noise with variances of 0.5/0.8 to assess model performance under severe visual distortion. As noise increases, all models struggle with geometric recognition, showing significant performance drops. Fig.10 reports perception accuracy under noise, and distorted examples are shown in Appendix Fig.11, where at variance 0.8, even humans fail to interpret the diagrams.
Table 5: Performance of relationship identification (rlat) on MATHGLANCE under different textual distractor settings.
MATHGLANCE (rlat) Unrela. infor. Conflicts
w/o w/Δ↑↓ w/o w/Δ↑↓
Qwen2VL-7B 53.0 48.5↓4.5 53.0 24.5↓28.5
Qwen2.5VL-7B 52.0 56.0↑4.0 52.0 30.0↓22.0
DeepSeek-VL2-Tiny 32.0 34.5↑2.5 32.0 3.5↓28.5
DeepSeek-VL2-Small 48.5 42.5↓6.0 48.5 20.5↓28.0
LLaVa-1.5-13B 42.0 37.5↓4.5 42.0 16.0↓26.0
InternVL2.5-8B 60.0 49.0↓11.0 60.0 18.0↓42.0
GPT-4o 62.5 63.5↑1.0 62.5 38.0↓24.5
SVE-Math-DeepSeek-7B 51.0 49.0↓2.0 51.0 15.0↓36.0
SVE-Math-DeepSeek +-7B 88.0 86.5↓1.5 88.0 74.0↓14.0
Text Distractors. Researchers have observed that large language models (LLMs) often dominate MLLMs, relying heavily on knowledge shortcuts rather than genuinely perceiving visual content. To investigate this, we evaluate two settings in the relationship identification task. The first involves introducing unrelated information w.r.t. the target task. For example, when the task is to identify a mathematical relationship, the model is additionally given irrelevant details about foreground/background colors, box shapes, or object attributes. The second setting introduces contradictory cues; for instance, if two lines are parallel, the model is explicitly prompted that they are perpendicular. The results in Tab. 5 show that providing conflicting knowledge significantly impairs the perceptual ability of all evaluated MLLMs. The impact is especially pronounced in models with weaker perception capabilities, such as InternVL2.5-8B (48.4% in avg.), while stronger perceptual models like Qwen2.5-VL-7B (59.2% in avg.) are less affected. Our method demonstrates greater robustness, though it still experiences a 14% performance drop. These findings highlight that reliance on knowledge shortcuts remains a critical issue, particularly in abstract diagram recognition, where models must genuinely understand visual content rather than rely on superficial pattern matching.
Chain-of-Thought (CoT) Response. CoT reasoning is designed to enhance step-by-step logical inference, yet its effectiveness in visual perception tasks remains uncertain. Our analysis shows that while CoT improves textual reasoning, it does not directly enhance spatial or geometric understanding. Models incorporating CoT reasoning often struggle with fundamental perception tasks, leading to significant CoT errors. Models often generate excessive yet irrelevant rationale misaligned with the diagram, ultimately resulting in incorrect responses, particularly in Qwen-series models and GPT-4o (over 30% in Fig. 6). See §G for CoT response examples in the Appendix.
Table 6: Performance comparison w.r.t. variants trained on GeoPeP and other mathematical visual alignment datasets. $\bigstar$ indicates the pretrained model continuously self-fine-tuned using MathV360K and instruction-formatted GeoPeP.
Model MATHGLANCE MathVerse MathVista
Plane Soild Graphs
SVE-Math-DeepSeek 35.4 49.4 55.1 24.3 48.7
Align. +AutoGeo ($\triangle$) 18.6 34.7 40.4 17.8 45.0
+MAVIS ($\Diamond$) 6.0 2.5 6.9 6.8 34.3
+GeoPeP ($\heartsuit$) 41.6 40.2 49.2 19.7 46.2
SFT $\triangle$ $\bullet$ $\bigstar$ 79.6 44.2 55.1 25.2 48.3
$\Diamond$ $\bullet$ $\bigstar$ 78.1 42.3 54.1 23.3 47.1
$\heartsuit$ $\bullet$ $\bigstar$ 84.6 54.1 60.7 28.1 51.3
3.3.2 Effect of GeoPeP
Building on SVE-Math-DeepSeek, we first train the projector while freezing the LLM and visual encoder during the visual-language alignment stage using either AutoGeo, MAVIS, or GeoPeP image-caption alignment datasets. As shown in Tab. 6, training with GeoPeP yields a 6.2% gain on perception tasks, but suffer a substantial drop on the other two datasets ($\sim$-27%). Using these three pretrained models, we then conduct self-fine-tuning (SFT) with MathV360K and instruction-formatted GeoPeP. The model trained with GeoPeP achieves a notable $\sim$4improvement on both MathVerse and MathVista compared to SVE-Math-DeepSeek, whereas other variants underperform on MathVista. We hope this observation encourages further research into more effective methods and training datasets to enhance the visual perception capabilities of mathematical MLLMs, as clearer visual understanding enables more faithful and accurate reasoning responses (Figs.7/15-16).
4 Conclusion
In this paper, we introduce MATHGLANCE, a perception-demanding benchmark designed to evaluate geometric perception in MLLMs. While these tasks appear trivial for humans to solve ‘at a glance’, they present significant challenges for current models. Unlike semantic-rich natural images, mathematical diagrams are structured and abstract, characterized by symbolic elements and their precise interrelationships. Thus, they require genuine visual interpretation rather than pattern recognition or textual shortcuts. Through extensive analysis, we identify key factors affecting diagam perception, highlighting the importance of high-quality and structure-aware training datasets to enable MLLMs to both see and reason effectively. Our findings suggest that improving low-level geometric perception is essential for advancing high-level reasoning, bridging the gap between visual understanding and logical inference.
Appendix
Appendix A Related Work
A.1 Mathematical Reasoning Benchmark
A.2 MLLMs for Math
Appendix B Visualization of MATHGLANCE
MATHGLANCE is specifically designed to evaluate perception-demanding tasks in Multimodal Large Language Models (MLLMs), focusing on four core visual understanding capabilities: shape classification, object counting, object grounding, and relationship identification. These tasks, commonly studied in classical computer vision, are typically solvable by humans with minimal cognitive effort. Our benchmark spans three key domains—plane geometry, solid geometry, and mathematical graphs—to ensure broad coverage of visual representations encountered in educational and scientific contexts. Fig. 8 provides illustrative examples from MATHGLANCE, demonstrating the diagrammatic input and corresponding question–answer (Q&A) pairs used for evaluation.
Figure 8: Visualization of sample cases from MATHGLANCE. (a), (b), and (c) correspond to problems related to Plane Geometry, Solid Geometry, and Graphs, respectively.
Appendix C Problem Templates
This section introduces the problem templates employed in MATHGLANCE, with illustrative examples provided in Tab. 7. We adopt a template-based generation engine to systematically construct diverse perception-oriented Q&A pairs. Each question is generated by parsing the structured JSON annotations of a given image, which include information on shapes, object attributes, spatial positions, and inter-object relationships. These elements are then filled into carefully designed templates (Tab. 7) to produce grammatically correct and semantically meaningful Q&A pairs.

### v2

@@ A.2 MLLMs for Math
Table 9: Key statistics of MATHEMETRIC are summarized in Tab. 7(a), and the subject-task distribution is illustrated in Fig. 7(b).
@@
Appendix B Detailed Data Statistics of MATHEMETRIC
To comprehensively evaluate the perceptual abilities of MLLMs in mathematical contexts, MATHEMETRIC includes 1,609 questions grounded in 1,198 unique images. These span three key domains: plane geometry (66%), solid geometry (20%), and graphical representations (14%), such as line, bar, and pie charts (see Tab. 7(b)). The benchmark comprises four core perception tasks: shape classification (489 questions, 30.4%), object counting (401, 24.9%), relationship identification (313, 19.5%), and fine-grained object grounding (406, 25.2%). Except for the grounding task, which uses free-form answers (e.g., bounding box coordinates), all other tasks are framed as multiple-choice (55.5%) or true/false (19.3%) formats to support scalable and consistent evaluation. By default, ground truth answer is randomly assigned with equal probability, 25% for each choice letter and 50% for true/false. In our benchmark: the proportion of answer A/B/C/D is 26.7%/21.8%/28.3%/23.2%; and 48.3%/51.7% for true/false. As summarized in Tab. 7(b), the dataset includes 1,514 unique questions and 380 unique answers, with an average question length of 118.0 tokens and answer length of 1.8 tokens. This design supports fine-grained analysis of model performance across diverse perceptual challenges and subject domains.
Appendix C Training Implementation Details
We conduct experiments on both the math-specific SVE-Math-DeepSeek, and generic Qwen2.5-VL models (7B and 32B). Both follow a two-stage training pipeline: i) alignment training with image-caption pairs and ii) instruction tuning for reasoning and perception.
For SVE-Math-DeepSeek, we first train the vision-language projector using image-caption pairs (GEOMETRIC) while keeping the visual encoder and language model frozen. In the second stage, we perform full-parameter self-fine-tuning (SFT) using a combination of instructive GEOMETRIC samples and MathV360K, updating the visual encoder, projector, and LLM jointly. For the generic Qwen2.5-VL-7B/32B models, we follow the same first-stage alignment process. To preserve prior reasoning capabilities and mitigate catastrophic forgetting, we adopt parameter-efficient LoRA tuning and incorporate an additional 300K MultiMath [23] visual reasoning samples in the second SFT stage.
Training Settings. During alignment training, we use a global batch size of 256, a learning rate of $1\mathrm{e}{-4}$, and train for one epoch with a maximum sequence length of 2048. In the SFT phase, we reduce the learning rate to $1\mathrm{e}{-5}$ (with a vision-specific learning rate of $2\mathrm{e}{-6}$) and maintain the same batch size and sequence length. We use the Adam optimizer without weight decay and apply a cosine learning rate schedule. To improve memory efficiency, we employ Fully Sharded Data Parallel (FSDP), gradient checkpointing, and enable BF16 precision. We avoid CPU/GPU offloading to maximize throughput.
Hardware and Runtime. All 7B model training was performed on 8 $\times$ A100 GPUs (80GB each). Alignment training took approximately 5 hours, followed by 12 hours for full fine-tuning (SVE-Math-DeepSeek) and 16 hours for LoRA fine-tuning (Qwen2.5-VL-7B). For the 32B Qwen2.5-VL model, we used 16 $\times$ H100 GPUs (96GB each), with alignment training taking 9 hours and LoRA-based SFT 15 hours.
More Ablations. 1) Herein, we assess whether it is necessary to count complex overlapping objects. Even humans interpret overlapping objects differently. To ensure more accurate conclusions for overlapping object counting, we asked three authors to manually and independently label the same 200 images sampled from our benchmark. The ground truth for these cases is represented as a list of one to three possible answers (if three annotators provided the same answer, we merged them). We added a system prompt to our counting templates: you are required to count potential overlapping objects. For evaluation, we adopt free-form answer matching—if the model’s response matches any entry in the annotated answer list, it is considered correct. Top-1 accuracy is then computed based on this criterion, as shown in Tab 10.
Table 10: Accuracy (%) on non-overlapping vs. overlapping counting. Overlapping objects remain consistently harder for MLLMs.
LLaVA-v1.5-7B LLaVA-v1.5-13B G-LLaVA-7B MultiMath-7B Qwen2.5VL-7B InternVL2.5-8B InternVL2.5-38B GPT-4o SVE-Math-DeepSeek-7B SVE-Math-DeepSeek +-7B
Non-overlapping 43.8 44.5 45.8 35.9 54.8 40.0 56.2 57.1 40.1 88.2
Overlapping 28.7 29.1 29.3 18.4 32.8 21.7 31.9 38.9 23.7 67.3
2) Instead of using numeric-coordinate grounding, we adopt vertex lists as an alternative abstraction for grounding planar geometric objects. Vertex-list grounding is conceptually simpler than coordinate-based approaches: rather than predicting multiple continuous numerical values, the model only needs to identify discrete symbolic labels already annotated in the diagram. For example, Q: Please provide the list of vertex labels described by the sentence: Square. A: ABCD. A comparison between vertex-list grounding and numeric-coordinate grounding is shown in Tab. 11.
Table 11: Comparison of vertex-list grounding vs. numeric-coordinate grounding for planar diagrams.
Qwen2-VL-7B Qwen2.5-VL-7B LLaVA-v1.5-7B InternVL2.5-32B GPT-4o G-LLaVA-7B MultiMath-7B SVE-Math-DeepSeek-7B SVE-Math-DeepSeek +-7B
Vertex 45.6 49.1 1.8 42.7 44.8 3.9 33.1 41.7 89.3
Numerics 12.8 18.5 14.2 2.5 1.1 0.4 1.1 3.6 82.9
3) To prevent models from exploiting vertex-count shortcuts in the shape-classification task, we design hard distractors for 75% of the questions. For each ground-truth polygon, at least one distractor is drawn from the same polygon family, ensuring that models cannot rely solely on the number of vertices. For instance, when the correct answer is a scalene triangle, distractors are sampled from other triangle types (isosceles, right, equilateral). Likewise, for quadrilaterals, distractors include shapes such as rectangles, squares, or trapezoids (e.g., a rectangle paired with a right trapezoid; see Fig. 9, first row). To further validate that vertex-count shortcuts are not driving model performance, we construct an additional controlled dataset of 200 planar diagrams in which vertex labels are replaced with purely visual numeric markers (e.g., 1, 2, 3 $\cdots$). These markers do not encode polygon type and therefore eliminate vertex-count cues. Example cases are provided in Fig. 8, and the corresponding ablation results are summarized in Tab 12. These results show diverse behavior across models when switching from vertex labels to visual markers. If vertex-count shortcuts dominated model performance, all models would exhibit similar, substantial drops under the marker condition—but this is not observed. Several models (e.g., Qwen2.5-VL-7B, LLaVA-v1.5-7B, SVE-Math-DeepSeek-7B, and our model) maintain stable performance under both conditions. The observed irregularities instead reflect differences in models’ ability to interpret visual markers: for example, GPT-4o handles marker-based diagrams well, whereas InternVL2.5-38B performs better with vertex labels. These findings confirm that our distractor design is robust and that model performance is not driven by vertex-count shortcuts.
Table 12: Accuracy (%) on vertex-list vs. visual-marker classification for planar diagrams.
Qwen2-VL-7B Qwen2.5-VL-7B LLaVA-v1.5-7B InternVL2.5-38B GPT-4o G-LLaVA-7B MultiMath-7B SVE-Math-DeepSeek-7B SVE-Math-DeepSeek +-7B
Vertex 50.7 48.6 34.6 55.3 53.1 31.6 44.1 43.9 72.6
Markers 43.3 47.9 35.3 43.3 54.8 28.8 33.2 41.1 71.7
Figure 8: Visualization of samples where objects are annotated using numerical visual markers instead of vertex labels. Zoom in for the best view.
Discussion of MINT-CoT and Vision-R1 vs. Ours. MINT-CoT and Vision-R1 align with our findings that perception is critical for reasoning, yet operate at different scopes:
1) Relation to MINT-CoT: MINT-CoT explicitly injects interleaved visual tokens into the chain-of-thought, enabling the model to reference visual evidence during reasoning. This represents an important direction for coupling perception and reasoning step-wisely. In contrast, our work focuses on improving diagram perception itself, without introducing any specialized mechanism for mixing visual and textual tokens in the reasoning process. As a result, the two lines of work are compatible: MINT-CoT studies how visual information is integrated into CoT reasoning; our work studies whether the model can accurately perceive symbolic diagrams in the first place, and how enhanced perception naturally transfers to reasoning.
2) Relation to Vision-R1: Vision-R1 aims to incentivize reasoning ability on top of a strong base model using reinforcement learning. A key factor in Vision-R1’s success is the quality and diversity of its cold-start data: over 43 curated datasets spanning mathematical diagrams, science and medical figures, general QA images, and figure-understanding datasets. Importantly, during data construction, Vision-R1 generates pseudo-CoT using image captions, giving the model holistic visual context that supports the development of high-quality reasoning traces. While both works support the idea that perception is a prerequisite for effective multimodal reasoning, Vision-R1 focuses primarily on improving reasoning ability, whereas our work focuses on explicitly enhancing symbolic diagram perception. This distinction is crucial: reasoning improvements in Vision-R1 rely heavily on RL strategies and large-scale, high-quality CoT training samples, not on explicit perceptual enhancement.
Overall, we summarize the key points below:
  • Our work: focuses on improving diagram perception and demonstrates measurable transfer from enhanced perception to downstream reasoning.
  • MINT-CoT: interleaves visual tokens within chain-of-thought reasoning to produce more visually aligned inference.
  • Vision-R1: incentivizes reasoning ability through reinforcement learning, built on high-quality, large-scale perception–reasoning training data.
Appendix D Visualization of MATHEMETRIC
MATHEMETRIC is specifically designed to evaluate perception-demanding tasks in Multimodal Large Language Models (MLLMs), focusing on four core visual understanding capabilities: shape classification, object counting, object grounding, and relationship identification. These tasks, commonly studied in classical computer vision, are typically solvable by humans with minimal cognitive effort. Our benchmark spans three key domains—plane geometry, solid geometry, and mathematical graphs—to ensure broad coverage of visual representations encountered in educational and scientific contexts. Fig. 9 provides illustrative examples from MATHEMETRIC, demonstrating the diagrammatic input and corresponding question–answer (Q&A) pairs used for evaluation.
Figure 9: Visualization of sample cases from MATHEMETRIC. (a), (b), and (c) correspond to problems related to Plane Geometry, Solid Geometry, and Graphs, respectively.
Appendix E Problem Templates
This section introduces the problem templates employed in MATHEMETRIC, with illustrative examples provided in Tab. 13. We adopt a template-based generation engine to systematically construct diverse perception-oriented Q&A pairs. Each question is generated by parsing the structured JSON annotations of a given image, which include information on shapes, object attributes, spatial positions, and inter-object relationships. These elements are then filled into carefully designed templates (Tab. 13) to produce grammatically correct and semantically meaningful Q&A pairs.